OpenStudy (anonymous):

find ∞ ∑ 50(2x-1) ^ r-1 r=1 when x=0.3

OpenStudy (amistre64):

the sum of a geometric series is:\[\frac{a}{1-r}\]

OpenStudy (amistre64):

in this case: a = 50(.6-1)

OpenStudy (amistre64):

.... and so does r I think

OpenStudy (amistre64):

50(-.4)^(r-1) a = 50 and r=-.4 i prolly shoulda used different letters so as not to confuse the original "r"

OpenStudy (anonymous):

hmmm ok... but i dont think this is completed yet

OpenStudy (amistre64):

since |-.4| < 1 this should converge

OpenStudy (amistre64):

\[\frac{50}{1--.4}=\frac{50}{1.4}\]